OpenStudy (anonymous):

A salesman drives from Ajax to Barrington, a distance of 120 mi, at a steady speed. He then increases his speed by 10 mi/h to drive the 150 mi from Barrington to Collins. If the second leg of his trip took 6 min more time than the first leg, how fast was he driving between Ajax and Barrington?

OpenStudy (anonymous):

Please help I'm confused!

OpenStudy (anonymous):

i am sure we can do this, let me write it down so i don't make a mistake

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

ok lets put the rate from A to B as \(r\) then since time is distance divided by rate , we can write \[t=\frac{120}{r}\] where \(t\) is the time from A to B

OpenStudy (anonymous):

the rate from B to C is 10 more, i.e. \(r+10\) and the time is 6 more, i.e. \(t+6\) so we also have \[t+6=\frac{150}{r+10}\]

OpenStudy (anonymous):

then the job is to solve the two equations \[t=\frac{120}{r}\\t+6=\frac{150}{r+10}\]

OpenStudy (anonymous):

we can get rid of the \(t\) in the second one and write \[\frac{120}{r}+6=\frac{150}{r+10}\]

OpenStudy (anonymous):

wait i think i screwed this up, hold the phone

OpenStudy (anonymous):

Oh man okay

OpenStudy (anonymous):

Now you know why I'm confused haha

OpenStudy (anonymous):

lol i am an idiot it is six MINUTES not 6 hours

OpenStudy (anonymous):

Lol exactly

OpenStudy (anonymous):

And you're not an idiot lol

OpenStudy (anonymous):

\[\frac{120}{r}+\frac{1}{10}=\frac{150}{r+10}\]

OpenStudy (anonymous):

as six minutes is one tenth of an hour

OpenStudy (anonymous):

What would be the unit for r?

OpenStudy (anonymous):

r is a rate, it is miles per hour t is also in hours i changed 6 minutes to one tenth of an hour

OpenStudy (anonymous):

so now all the units match up , miles, hours, miles per hour

OpenStudy (anonymous):

Do you mind solving the equation, because this is actually a question from the practice test and I'm a visual learner and for me to comprehend it more clearly ill need all the help I can get please!

OpenStudy (anonymous):

i guess we can do it \[\frac{120}{r}+\frac{1}{10}=\frac{1200+r}{10r}\]

OpenStudy (anonymous):

then we get \[\frac{1200+r}{10r}=\frac{150}{r+10}\] "cross multiply" and get \[(r+10)(1200+r)=1500r\]

OpenStudy (anonymous):

multiply out, set it equal to zero and solve you get \[r^2-290r+12000=0\] and by some miracle this factors as \[(r-50)(r-240)=0\]

OpenStudy (anonymous):

so either \(r=50\) or \(r=240\) i think we should pick 50

OpenStudy (anonymous):

then we still have to find the time from A to B but now that is easy, as it is distance divided by rate, which is \(\frac{120}{50}=2.4\) or two hours and 24 minutes

OpenStudy (anonymous):

i have to say this was a pain, and there may have been a shorter way to do it, but i don't see it

OpenStudy (anonymous):

Lol it's okay but thanks! :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

So how would I know I'll have to put the distance over r?